OpenStudy (anonymous):

A part-time landscaper made $8996.32 last year. If she claimed herself as an exemption for $3650 and had a $5700 standard deduction, what was her taxable income last year? A. $0 B. $5346.32 C. $353.68 D. $3296.32

OpenStudy (anonymous):

what school do you go to

OpenStudy (anonymous):

????

OpenStudy (anonymous):

dingdong high

OpenStudy (texaschic101):

Exemption means not subject to taxation. Deduction means taking some amount of your income for the year, and not have to pay taxes on it. 8996.32 - (3650 + 5700) = 8996.32 - 9350 = - 353.68 ANSWER : A She does not have to pay. Her taxable income is below $0.